A pair of shoes cost 32cedi more than a football if the cost of the pair of shoes is 95 cedis what is the cost of the football


1
Expert's answer
2021-07-15T11:23:30-0400

We proceed if we consider S = price of a pair of shoes and F = price of a football:


"S = F + 32\,cedis
\\S = 95 \,cedis
\\\therefore F=S-32\,cedis 
\\ \implies F= 95 \,cedis- 32\,cedis=63\,cedis"


In conclusion, a football has a cost of 63 cedis.